# Sparse is Enough in Fine-tuning Pre-trained Large Language Model

### v1

@@ Abstract
With the prevalence of pre-training-fine-tuning paradigm, how to efficiently adapt the pre-trained model to the downstream tasks has been an intriguing issue. Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed for low-cost adaptation, including Adapters, Bia-only, and the recently widely used Low-Rank Adaptation. Although these methods have demonstrated their effectiveness to some extent and have been widely applied, the underlying principles are still unclear. In this paper, we reveal the transition of loss landscape in the downstream domain from random initialization to pre-trained initialization, that is, from low-amplitude oscillation to high-amplitude oscillation. The parameter gradients exhibit a property akin to sparsity, where a small fraction of components dominate the total gradient norm, for instance, 1% of the components account for 99% of the gradient. This property ensures that the pre-trained model can easily find a flat minimizer which guarantees the model’s ability to generalize even with a low number of trainable parameters. Based on this, we propose a gradient-based sparse fine-tuning algorithm, named Sparse Increment Fine-Tuning (SIFT), and validate its effectiveness on a range of tasks including the GLUE Benchmark and Instruction-tuning. The code is accessible at https://github.com/song-wx/SIFT.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across various tasks. With the prevalent pre-training-fine-tuning paradigm, LLMs can be effectively adapted to the specific downstream domain. As the sizes of pre-trained language models (PLMs) scale up, e.g. GPT-3(Brown et al., 2020) with 175 billion parameters, fine-tuning all the parameters (full fine-tune) has been prohibitively expensive. Therefore, toward efficient adaptation from PLMs to the downstream tasks, several parameter-efficient fine-tuning methods have been proposed, which can match the performance close to or even surpass that of full fine-tuning with a relatively low costs.
@@
Although these methods have demonstrated their effectiveness to some extent and have been widely applied, the underlying principles are still unclear. These methods differ in terms of the selection of trainable parameters and the structure of the trainable modules, but the essence is to train a small set of parameters in order to lead the pre-trained model towards a relatively optimal minimum in downstream domains. This is consistent with (Li et al., 2018a) that over-parameterized networks have a lower intrinsic dimension, i.e. they can achieve 90% or even 100% of the full parameter performance by training only fewer parameters than the full parameter, which is the source of LoRA’s motivation. This has also been verified by numerous PEFT methods, where just fine-tuning a fairly small parameter (e.g. 0.5% of the full parameter) can allow the pre-trained model to be effectively adapted to downstream tasks.
Even though we recognize that the intrinsic dimension of the over-parameterized model is small, in intuition, it is not straightforward to select the correct, appropriate subspace to represent the information of the full space. The existing PEFT methods all go about constructing the subspace heuristically. A unified perspective on the mainstream PEFT methods is given in (He et al., 2022) which argues that from a functional perspective, these methods can all be seen as the form of $down\ projection\to non-linearity\to up\ projection$. To this end, a natural question is:
@@
What is the essence of the low intrinsic dimension of pre-trained language models?
To answer this question, we need to understand what happens during the transition from a randomly initialized model to a pre-trained model. This paper investigates the structure of the loss landscape in downstream domains, viewing the essence of pre-training as guiding the model from a random initialization to a location in the loss landscape with a distinctive structure. We gain an intuitive understanding of this change by visualizing the loss landscape from random parameters to pre-trained parameters. Through one-dimensional and two-dimensional visualizations of the loss landscape, Figure 1 demonstrates a general property: by pretraining, the loss landscape transforms from low-amplitude oscillations near random initialization to high-amplitude oscillations. Corroborating this property, we find that pre-trained models exhibit a sparse-like characteristic in their parameter gradients for downstream tasks, where a small fraction of components account for the majority of the total gradient norm. For example in Figure 2, 1% of the components constitute 99% of the gradient. Hence, we attribute the low intrinsic dimension to this sparse-like property of the gradient space and demonstrate that LoRA’s low-rank structure effectively captures this characteristic. Based on this principle, we propose a gradient-based sparse update algorithm, which involves adding hooks during the backward propagation to acquire a portion of the gradients, thereby achieving both parameter and memory efficient during fine-tuning. We validate its effectiveness on a range of tasks including the GLUE Benchmark and Instruction-tuning.
@@
  • To our best knowledge, this is the first time revealing a gradient characteristic widely present in pre-trained models on downstream tasks, where a small fraction of components account for the majority of the gradient norm.
  • We propose a gradient-based sparse fine-tuning algorithm, an efficient and simple parameter-efficient fine-tuning method. This approach discards the need for well-designed structures common in various previous PEFT methods, relying only on gradient information to estimate the importance of components and achieving efficient fine-tuning by updating only a part of the components.
  • We introduce a memory-efficient sparse fine-tuning approach by utilizing hooks during backward propagation, which could potentially be implemented at the lower level of deep learning training frameworks like PyTorch(Paszke et al., 2017) in the future. This paves the way for more streamlined and efficient fine-tuning, holding considerable potential for widespread application.
@@ 2 Related Work
Pre-training-fine-tuning Paradigm: Pre-training models are extensively used in several fields of artificial intelligence, including natural language processing, computer vision and graph learning. Pre-training models learn low-level knowledge by training on large-scale data and thus, from an initialization point of view, pre-training provides better initialization parameters to various downstream tasks. In the field of Natural Language Processing(NLP), language models(LMs) are pre-trained on a large corpus in order to learn rich syntactic and semantic knowledge. There are two main research directions in pre-trained language models (Zhou et al., 2023), one is the autoencoder LMs represented by BERT (Devlin et al., 2019), which learns by training the bi-directional encoders in the Transformer to predict masked words and determine the contextual relationships of sentences. Based on BERT, (Liu et al., 2019) optimized the pre-training method, and in this paper, RoBERTa is used as the backbone model to validate the effectiveness of Sparse Increment Fine-Tuning(SIFT) in the GLUE benchmark.The other is the autoregressive LMs represented by GPT (Radford & Narasimhan, 2018), which extracts the features through autoregressive decoders, and based on the previous words to predict the next word, and fine-tuning on downstream data to solve a specific task. Currently, the majority of the mainstream large language models adopt the autoregressive model with decoder-only, such as Llama (Touvron et al., 2023). In order to allow the LMs to follow human intentions and produce more meaningful content, (Wei et al., 2022) proposes to fine-tune the language model with instructions after pre-training, enabling LMs to align instructions. This paper tests the performance of pre-trained Llama using SIFT for Instruction-tuning.
Figure 1: 1-D (up) and 2-D (down) visualization of the transition of the loss landscape from random initialization to pre-trained initialization. The figures show the loss landscape transitions from low amplitude oscillations to high amplitude oscillations.
@@
Parameter-efficient Fine-tuning: Although the pre-training-fine-tuning paradigm is becoming increasingly popular, the expense of fully fine-tuning large pre-trained models is prohibitively high. (Ding et al., 2022) points out that only 0.5-4% of research papers experimentally adopt large pre-trained models, due to the unaffordable costs of deploying and fine-tuning these large models. To apply pre-trained models more efficiently to downstream tasks, (Houlsby et al., 2019) proposed parameter-efficient transfer learning by inserting a bottleneck structure between different layers of the pre-trained model. During fine-tuning, the pre-trained parameters remain unchanged, and only the bottleneck layers are updated, achieving parameter-efficient fine-tuning. (Zaken et al., 2022) shows that training only the bias terms and task-specific parts of the model during the fine-tuning process can also adapt the pre-trained models effectively to downstream tasks. Prefix-tuning (Li & Liang, 2021) adapts the pre-trained model to downstream tasks by prepending trainable continuous tokens (Prefixes) to the input and hidden states of each Transformer layer. LoRA (Hu et al., 2021) freezes the weights of the pre-trained model and injects trainable rank decomposition matrices into each layer of the Transformer architecture. Following the proposal of LoRA, several LoRA-based improvements have been introduced, including AdaLoRA (Zhang et al., 2023), which adaptively allocates parameter budgets based on the importance of weight matrices, and QLoRA (Dettmers et al., 2023), which further reduces the resource overhead for fine-tuning large pre-trained models through quantization techniques. The above PEFT methods, except for Bias-only which is a module sparse approach, are all adapted to the downstream tasks by introducing additional modules with fewer parameters. In this paper, we present a component sparse approach to achieve almost full fine-tuning performance without introducing additional parameters.
Loss Landscape : Loss landscape is a high-dimensional geometry which is completely determined as soon as a dateset and the network structure are specified(Li et al., 2018a). Our intuition may not be reliable in higher dimension and it is not uncommon for deep learning models of loss landscapes to hold special properties. (Hochreiter & Schmidhuber, 1997) (Keskar et al., 2017) point out that the flatter minima in loss landscape found by stochastic gradient based methods results in better generalization. For the pre-trained models, the flatness of the models on the loss landscape strongly correlates to the performance of the downstream task, even if the pre-training losses are the same (Liu et al., 2022). One approach for analyzing the loss landscape is to visualize this high-dimensional geometry in one or two dimensions. (Li et al., 2018b) visualizes the loss landscape by introducing a ”filter normalization” method, and explains the role of skip connection in terms of the (non)convexity of the loss landscape. (Hao et al., 2019) visualizes the trajectory in the loss landscape during fine-tuning BERT for downstream task and compares the difference between the trajectory of training from scratch and of training from pre-trained model. It is shown that pre-training enables the loss to converge quickly to flatter minima. In this paper, we visualize the transition in loss landscape between random initialization and pre-trained initialization based on the method of (Hao et al., 2019; Li et al., 2018b).
@@
3 Understanding Pre-training-fine-tuning from a Gradient-based Perspective
The pre-training-fine-tuning paradigm has shown great success in numerous NLP tasks (Devlin et al., 2019), to our best knowledge, there is still no conclusive answer as to why pre-training can improve performance and generalization ability in different tasks. In this section we will try to understand the pre-training fine-tuning paradigm from the perspective of gradient distribution along with the visualization of loss landscape.

### v3

@@ Abstract
With the prevalence of pre-training-fine-tuning paradigm, how to efficiently adapt the pre-trained model to the downstream tasks has been an intriguing issue. Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed for low-cost adaptation. Although PEFT has demonstrated effectiveness and been widely applied, the underlying principles are still unclear. In this paper, we adopt the PAC-Bayesian generalization error bound, viewing pre-training as a shift of prior distribution which leads to a tighter bound for generalization error. We validate this shift from the perspectives of oscillations in the loss landscape and the quasi-sparsity in gradient distribution. Based on this, we propose a gradient-based sparse fine-tuning algorithm, named Sparse Increment Fine-Tuning (SIFT), and validate its effectiveness on a range of tasks including the GLUE Benchmark and Instruction-tuning. The code is accessible at https://github.com/song-wx/SIFT.
Parameter-Efficient Fine-Tuning, Deep Learning, Machine Learning
@@ 1 Introduction
With the prevalent pre-training-fine-tuning paradigm, Large Language Models (LLMs) pre-trained on a internet-scale corpus can be effectively adapted to the specific downstream domains. However, as the number of parameters in large language models increases, fine-tuning all parameters is prohibitively expensive. Toward efficient adaptation from pre-trained LLMs to downstream tasks, several parameter-efficient fine-tuning methods have been proposed (Houlsby et al., 2019; Ben Zaken et al., 2022; Li & Liang, 2021; Hu et al., 2021; Zhang et al., 2023; Dettmers et al., 2024), which can achieve the performance close to or even surpass that of full fine-tuning with lower costs.
@@
Although these methods have demonstrated their effectiveness to some extent and have been widely applied, the underlying principles are still unclear and the selection of trainable modules are typically heuristically. Li et al. (2018a) points out that over-parameterized networks have a lower intrinsic dimension, i.e. they can achieve 90% or even 100% of the full parameter performance by training only fewer parameters than the full parameter, which is the motivation of LoRA (Hu et al., 2021). A natural question is:
@@
What causes pre-trained language models to have the lower intrinsic dimension?
To answer this question, we first need to understand what changes occur from random initialization to pre-trained initialization. As indicated in Devlin et al. (2018), pre-trained models rapidly generalize to downstream data through fine-tuning. Intuitively, the generalization error of pre-trained models is controlled by a tighter bound compared to random initialized models. Therefore, we shed a spotlight on a classic PAC-Bayesian generalization error bound to analysis the difference between training from scratch and training after pre-training, which is first proposed by McAllester (2003) and refined by Maurer (2004); Catoni (2007); Thiemann et al. (2017). In contrast to general PAC bounds that only consider the complexity of the hypothesis space, typically measured by VC-dimension (Vapnik, 1991) or Rademacher complexity (Bartlett & Mendelson, 2002), PAC-Bayesian bounds jointly consider a less accurate prior probability distribution for parameters and a more precise data-based posterior distribution. We view pre-training as a shift of prior, where pre-training, by learning language features such as grammar and syntax on a large corpus, moves the model away from sub-manifolds of non-expressive parameters in the parameter space. Pre-training initialization is equivalent to assigning relatively low probabilities in the prior for non-expressive parameters. In contrast to random initialization, which assigns equal probabilities in the prior to all parameters, the KL divergence between the pre-trained prior and the data-based posterior is lower, as shown in Figure 1. Therefore, pre-trained models have a tighter generalization error bound. We validate this prior shift through investigating changes in the loss landscape and gradient distribution. From random initialization to pre-trained initialization, the loss landscape transitions from flat oscillations to sharp oscillations, and the gradient exhibits a quasi-sparse property, where partial components dominate the majority of the gradient norm. Due to this difference, we conclude that the intrinsic dimensions actually reflects the required dimensions of the search space. For pre-trained models, the searching space dimensions needed to reach an optimal solution during the fine-tuning process has been compressed and the subspace spanned by the dimensions with extreme gradients largely contains the optima. As a result, the pre-trained models can be effectively adapted to downstream tasks by updating only a small portion of parameters.
Based on this, we propose a gradient-based sparse update scheme called Sparse Increment Fine-Tuning (SIFT). Unlike other existing PEFT methods, such as LoRA (Hu et al., 2021), Adapter (Houlsby et al., 2019), Prefix-Tuning (Li & Liang, 2021), which require pluging in additional modules, SIFT does not alter the model’s initial structure and is the first to achieve parameter efficiency in a component-sparse way. To achieve memory efficiency, it inserts hook functions into back propagation to acquire sparse gradients and perform in-place sparse updates on the parameters, as illustrated in Figure 5. We validate its effectiveness on a range of tasks including the GLUE Benchmark and Instruction-tuning and highlight its advantages in parameter efficiency in Section 5.3.2. Overall, we demonstrates the feasibility of fine-tuning language models entirely in a component-sparse way.
@@
  • To our best knowledge, this is the first time analyzing pre-trained models with PAC-Bayesian bounds. We validate the prior shift of pre-trained models from the perspectives of loss landscapes and gradient distributions and relate the low intrinsic dimensions to the compressed searching space.
  • We propose a gradient-based sparse fine-tuning algorithm, an efficient and simple PEFT method, which is an entirely component-sparse method that needs no additional plug-in modules. It demonstrates competitive performances and better parameter utilization efficiency.
  • We introduce a memory-efficient implementation for SIFT by utilizing hooks during backward propagation, which could potentially be implemented at the lower level of deep learning training frameworks like PyTorch (Paszke et al., 2017), and integrated into the general training process. This paves the way for more streamlined and efficient fine-tuning.
@@ 2 Related Work
Parameter-efficient Fine-tuning: To apply pre-trained models more efficiently to downstream tasks, Houlsby et al. (2019) proposes parameter-efficient transfer learning by inserting a bottleneck structure between different layers of the pre-trained model. During fine-tuning, the pre-trained parameters remain unchanged, and only the bottleneck layers are updated. Ben Zaken et al. (2022) shows that training only bias terms and task-specific modules during fine-tuning can also adapt pre-trained models effectively to downstream tasks. Prefix-tuning (Li & Liang, 2021) adapts pre-trained models to downstream tasks by prepending trainable continuous tokens (Prefixes) to the input and hidden states of each Transformer layer. LoRA (Hu et al., 2021) freezes the weights of pre-trained models and injects trainable rank decomposition matrices into each layer of the Transformer architecture. Following the proposal of LoRA, several LoRA-based improvements have been introduced, including AdaLoRA (Zhang et al., 2023), which adaptively allocates parameter budgets based on the importance of weight matrices, and QLoRA (Dettmers et al., 2024), which further reduces the resource overhead through quantization techniques. The above PEFT methods, except for Bias-only which is a module sparse approach, are all adapted to the downstream tasks by introducing additional modules.
PAC-Bayesian Generalization Error Bounds: introduced by McAllester (1999, 2003), combines the advantage of Bayesian methods and PAC Learning, which allows for considering domain knowledge as a Bayesian prior without assuming the truth of the prior (McAllester, 2003). Early generalization error bounds only consider the complexity of the hypothesis space, measured by VC-dimension (Vapnik, 1991) and Rademacher complexity (Bartlett & Mendelson, 2002), which is nearly vacuous and trivial in the real neural network setting (Zhang et al., 2021). In contrast, PAC-Bayes methods reward models whose prior assigns higher probability to those hypotheses that are in line with train data, with a tighter control over generalization error.
@@
3 Understanding Pre-training-fine-tuning from a Distribution-shift Perspective
The pre-training-fine-tuning paradigm has shown great success in numerous NLP tasks (Devlin et al., 2018). In this section, we adopt the PAC-Bayesian generalization error bounds and, from the perspective of the prior distribution shift, analyze the pre-training-fine-tuning paradigm in both theoretical and empirical aspects. We conclude that the low intrinsic dimensions of pre-trained models originate from the compressed searching space and the subspace spanned by the dimensions with extreme gradients largely contains an optima.
3.1 PAC-Bayesian Generalization Error Bounds
For a fixed architecture of the deep learning model, models with different parameter values (i.e., hypotheses) collectively form the hypothesis space. Any hypothesis $h$ in the hypothesis space $\mathcal{H}$ can be represented as $h=\sum_{i=1}^{N}h_{i}e_{i}$, where $N$ is the total number of parameters, $e_{i}$ is akin to the definition of the standard basis (with the $i^{th}$ component being 1 and the rest being 0), and $h_{i}$ represents the value of the model $h$ on the $i^{th}$ component. Hence, a straightforward conclusion is that $|\mathcal{H}|\leq N$, and subsequent discussions are based on the premise that the hypothesis space is finite.
Let the empirical risk of $h$ be ^$\hat{\mathcal{R}}(h)=\frac{1}{n}\sum_{i=1}^{n}l(h,s_{i})$, and the population risk be $\mathcal{R}(h)=\underset{s\sim\mathcal{S}}{\mathbb{E}}[l(h,s)]$, where $s_{i}$ is independently sampled from the identical data distribution $\mathcal{S}$, $n$ represents the number of samples, $l$ is a bounded loss function, and for simplicity, $l(h,s)\in[0,1]$. The generalization error is defined as $\Delta(h)=\hat{\mathcal{R}}(h)-\mathcal{R}(h)$.
In this setting, with a probability at least $1-\delta$, the population risk ^$\hat{\mathcal{R}}(h)$ satisfies:
$\forall h\in\mathcal{H},\hat{\mathcal{R}}(h)\leq\mathcal{R}(h)+\sqrt{\frac{log(|\mathcal{H}|)+log(1/\delta)}{2n}}$ (1)
If we consider a probability distribution $\mathcal{P}$ regarding hypotheses, referred to as a prior (McAllester, 1998), then with a probability of at least $1-\delta$,
$\forall h\in\mathcal{H},\hat{\mathcal{R}}(h)\leq\mathcal{R}(h)+\sqrt{\frac{\log(1/\mathcal{P}(h))+\log(1/\delta)}{2n}}$ (2)
A most trivial prior is to assume that all hypotheses are equally likely, i.e., $\forall h\in\mathcal{H},\mathcal{P}(h)=1/\log(|\mathcal{H}|)$. In this case, the inequality 2 degenerates to 1. But the parameters are certainly not uniformly distributed. For natural language processing tasks, without considering specific data, parameters that can accurately represent fundamental language features such as grammar and semantics should be assigned a higher probability. We refer to this as a weakly informative prior, which conveys partial information.
If we consider a more precise distribution $\mathcal{Q}$, which characterizes the probability distribution of the parameters based on specific data, referred to as a posterior, then we have the following bound on the population risk, first introduced by McAllester (2003), with a probability at least $1-\delta$:
$\underset{h\sim\mathcal{Q}}{\mathbb{E}}[R(h)]\leq\underset{h\sim\mathcal{Q}}{\mathbb{E}}[\hat{R}(h)]+\sqrt{\frac{\mathbb{KL}(Q||P)+\log({n/\delta})+2}{2n-1}}$ (3)
The formula 3 indicates that the smaller KL divergence between the prior and the posterior, the tighter control there is on the generalization error. A more detailed introduction of the above theorems can be found in Appendix A.
Pre-training language models aims to learn grammar, semantics, and other basic language features (referred to as a prior) from a large corpus through masking or auto-regressive methods, steering the model away from parameters that cannot correctly understanding and convey language. Therefore, pre-trained initialization of the model is equivalent to assigning lower probabilities to such parameters in the prior. Those parameters are naturally incapable of handling various tasks related to language correctly, so they should also have lower probabilities in the data-based posterior. As shown intuitively in Figure 1, pre-trained initialization considering basic language features is closer to the distribution of the data-based on posterior compared to random initialization. This closeness results in lower KL divergence and, consequently, a tighter upper bound on generalization error, allowing the population risk to consistently decrease along with the reduction of empirical risk during training.
The difference in the distribution of parameters can significantly affect the variation of losses, i.e., the properties of loss landscapes. In Section 3.2, we validate this distribution difference between random and pre-trained initialization by visualizing the loss landscape. By the first-order approximation of the multivariate Taylor expansion, this difference in loss landscapes stems from varied gradient distributions, which is also confirmed by the empirical analysis in Section 3.3.
Figure 1: An intuitive explanation about prior distribution and posterior distribution. Random initialization is equivalent to assigning equal probabilities to hypotheses in the hypothesis space (the horizontal line referred as prior-random). Pretraining learns language features from extensive corpora, moving away from hypotheses that are not accurately expressive of language, equivalent to assigning them lower prior probabilities. Data-based, more precise posterior also have lower probabilities for hypotheses that fail to represent and understand language correctly. Therefore, compared to prior-random, the KL divergence between prior-pretrained and posterior is smaller.
Figure 2: 1-D (up) and 2-D (down) visualization of the shift of the loss landscape from random initialization to pre-trained initialization. The figures show the loss landscape transitions from low amplitude oscillations to high amplitude oscillations.
3.2 Visualization of Loss Landscape
Loss landscape is a high-dimensional geometry which is completely determined as soon as a dateset and the network structure are specified (Li et al., 2018a), reflecting the relationship between the loss and parameters. For high-dimensional geometries such as loss landscapes, it is not possible to imagine and understand their geometric properties as intuitively as in low-dimensional shapes. Despite the difficulty in understanding such high-dimensional geometries, researchers have proposed methods to project high-dimensional loss surfaces into low-dimensional space, including one-dimensional visualization (Goodfellow et al., 2014; Smith & Topin, 2017) and two-dimensional visualization (Li et al., 2018b), in order to easily understand and study some of the properties of loss landscapes.
In exploring the shift of the loss landscape from random initialization to pre-trained initialization, we note $\theta_{0}$ as the randomly initialized parameters and $\theta_{1}$ as the pre-trained parameters. Goodfellow et al. (2014) plots the loss landscape curve from $\theta_{0}$ to $\theta_{1}$ by linearly sampling and interpolating the network parameters between $\theta_{0}$ and $\theta_{1}$, the loss curve $f$ is defined as:
$f(\alpha)$ $=\mathcal{L}(\theta_{0}+\alpha\delta)$
where $\alpha\in[0,1]$ and $\delta$ is $\theta_{1}-\theta_{0}$ that represents a direction. The two-dimensional loss surface $f$ is plotted similarly to the one-dimensional one, determined by a starting point and two directions:
In our setting, $\delta_{1}=\theta_{1}-\theta_{0},\alpha\in[-0.5,1.5],\beta\in[-1,1]$, when $\alpha=0,\beta=0$, the model is in the random initialization state, and when $\alpha=1,\beta=0$, the model is in the pre-trained state. We use a method similar to Li et al. (2018b) to construct the direction $\delta_{2}$ to avoid the scaling effect between different components, multiplying $\delta_{1}$ by a Gaussian random number points by points, i.e. $\delta_{2_{i}}=d_{i}*\delta_{1_{i}},$ where $d_{i}$ follows a normal distribution. We investigate the shift of the loss landscape using RoBERTa-base (Liu et al., 2019) on a range of tasks at GLUE, which are shown in Figure 2.
A prior idea would be that pre-trained models are more efficient at adapting to downstream tasks than training from scratch because the knowledge learned during pre-training reduces the initial loss. However, a striking fact is that pre-training does not make all tasks have lower losses than random initialization at the beginning of fine-tuning, e.g. MNLI, MRPC, etc. Instead, for most task, there are extremely higher losses than random initialization around the pre-trained initialization. A common feature of the loss landscape is that the oscillations of the loss around random initialization are generally small in magnitude, whereas the loss oscillations around pre-training are sharp, i.e. after a valid pre-training process, the loss landscape shifts from low amplitude oscillations to high amplitude oscillations.
Since the 1-D and 2-D loss landscapes are projections of the original loss landscape, the special states exhibited in the low-dimensional projections reflect the general property of the high-dimensional space. For simplicity, let’s consider the one-dimensional loss curve, $L(\theta_{0}+\alpha\delta)$, where $\delta$ can be represented as a linear combination of natural bases, i.e. $\delta=\sum_{i=1}^{N}d_{i}e_{i}$. By the first-order approximation of the multivariate Taylor expansion, we can estimate
$L(\theta_{0}+\alpha\delta)$ $=L(\theta_{0}+\alpha\sum d_{i}e_{i})$
$\approx L(\theta_{0})+\alpha\sum d_{i}\nabla L_{e_{i}}(\theta_{0})$